# A Survey of Optimization-based Task and Motion Planning: From Classical To Learning Approaches

### v3

@@ I Introduction
The main research focus in TAMP is to develop appropriate problem representations and algorithms that efficiently synthesize both symbolic and continuous components of the planning problem [2]. In the existing literature, there are three mainstream classes of TAMP methods: (i) constraint-based TAMP [4, 5], (ii) sampling-based TAMP [6, 7], and (iii) optimization-based TAMP [8, 9]. Constraint- and sampling-based TAMP characterizes the problem as a set of goal conditions. The solutions are typically found via constraint satisfaction or sampling-based approaches [10], which satisfy the defined goal conditions, but often cannot evaluate or compare the quality of the generated plan or the final state due to the lack of objective functions. In many robotics problems, goals are often expressed as an objective function rather than an explicitly defined set of states. For example, “Given a number of rectangular blocks on the table, build a stable structure that is as tall as possible with minimal robot control effort.” This is challenging for traditional sampling-based methods, which often require explicit goal definition and do not have mechanisms to compare plan qualities. As an exception, recent advancements in sampling-based motion planning [11] have been proposed to address optimal planning using $\text{RRT}^{*}$ and $\text{PRM}^{*}$ [12, 13]. For clarification, the scope of this survey focuses on optimization-based TAMP, which naturally defines an objective function for representing the plan quality, in addition to task- and motion-level constraints. This framework enables us to represent and solve a broad range of tasks with complex objective functions.
@@
However, the hybrid optimization problems formed by optimization-based TAMP are often computationally intractable. A successful planning algorithm needs to simultaneously overcome the combinatorial complexity at the task planning level, and the numerical complexity at the motion planning level. As such, a common theme in optimization-based TAMP is to trade off between the complexity of the optimization and comprehensiveness of the information included in the planning problem. Either extreme of this trade-off tends to degrade either the quality or the computational efficiency of the resulting robot plans. Therefore, the challenge remains to bridge the gap between planning for long-horizon tasks [21, 22], and generating highly dynamic robot behaviors, showcased in model-based optimal control strategies [23, 24, 25].
@@ I-A Survey Goals and Roadmap
This work is inspired by and builds upon previous surveys in TAMP [2, 10, 34, 35], but carries the unique overarching goal of reviewing the historical background and state-of-the-art optimization-based TAMP, and illustrating the connection between classical methods and the recent development in learning methods. Portions of this work are inspired by recent surveys in other relevant areas such as logic programming [22], formal methods [36], distributed optimization [37], and TO for legged locomotion [23].
@@ IV-B Learning Models for Task Planning
More recently, deep learning techniques have been explored to learn the models from large-scale datasets. Ames et al. [82] propose to learn preconditions, action parameters, and effects from execution results of parameterized motor skills, which are then used to construct symbolic models for efficient planning. Neural task programming [83] proposes to learn neural models that recursively decompose a task demonstration video into robot executable action primitives. To further improve the generalization on long-horizon tasks, Neural task graphs [84] learns neural networks for generating conjugate task graphs, where the actions are represented as nodes and the dependencies between actions are modeled by edges, better exploring the compositionality. egression planning networks [85] learns a neural model to iteratively predict the intermediate subgoals in a reverse order based on the current image observation and the final symbolic goal. Ceola et al. [86] propose to utilize deep reinforcement learning to train neural models for generating discrete actions. Deep affordance foresight [87] learns the long-term affordance of actions and the latent transition models to guide the search, and thereby informs the robot of the best actions to achieve the final task goal. Similarly, Liang et al. [88] propose to learn skill effect models that generate future terminal states of each parameterized skill, and then leverage these models to aid search-based task planning.
@@ IV-C LLMs for Task Planning
Traditionally, optimizing task plans for robots involves minimizing either the number of actions or the total plan cost, depending on whether action costs are considered. The emergence of LLMs, such as Google’s Bard, OpenAI’s ChatGPT [89], and Meta’s LLaMA [90], have reshaped the landscape of AI, including task planning for robots [91]. We categorize the LLM-based planning methods into the two groups: LLM-Native Planning Methods and LLM-Aided Planning Methods, where the former does not rely on external knowledge and the latter does, as shown in Fig. 7. While LLMs are not strong in numerical reasoning (and hence optimization) [92, 93], the incorporation of LLMs improves the capabilities of natural language understanding, the acquisition of world knowledge, and commonsense reasoning. Such capabilities make LLM-based task planners more friendly to end users, and better deal with tasks in open-world scenarios. LLM-Native planning methods often incorporate additional components like reinforcement learning to enhance planning by choosing better actions. Conversely, LLM-Aided planning methods can be integrated with classical optimization strategies, ensuring satisfactory planning efficiency and practicality. These two approaches are compatible with optimization methods, while integrating LLMs enhances the overall planning capabilities.
@@ V-B Distributed Optimization
Spatial Structure: The spatial structure of the system can be exploited when the subsystems and their dynamics are separable. This property often exists in multi-robot systems, where the planning of each robot can be treated as a sub-problem. The decision variables are possibily coupled through the objective functions or the constraints (e.g. collision avoidance between robots). Local copies of the full state variables can be created for each robot to decouple the problem [118], as seen in Fig. 8(a). Readers are referred to [37, 119] for a detailed review of the multi-robot ADMM. Robustness is further studied in [120] given a multi-robot motion planning problem via ADMM.
@@ VI Integrated Task and Motion Planning
The optimization-based formulations for TAMP often involve a hybrid optimization of discrete symbolic-level decisions and continuous motion-level trajectories, as shown in Sec. II-B. To this end, we identified two general approaches to formulate and solve the hybrid optimization: logic-guided TAMP and TO-guided TAMP. While both approaches inherently solve the hybrid optimization problems, they fundamentally differ in formulations and algorithms, particularly in the definition of discrete variables and the selection of search spaces. Fig. 9 shows the overall organization of this section and main topics discussed. Table I presents a representative set of classical approaches for optimization-based TAMP, highlighting their formulations, algorithms, and whether dynamics or kinematics are considered in the application.
@@
Table I: Overview of classical approaches for optimization-based TAMP
@@ VI-B3 Smooth Approximation Methods
Different from the hierarchical and distributed methods above, another line of research to circumvent the combinatorial complexity of MIP is reformulate the hybrid optimization into a continuous NLP with a specific cost function or constraint representing the smooth approximation of the discrete task planning (Fig. 12(b)). Such a formulation can be solved more efficiently with gradient-based solvers. Recent works in STL utilize smooth approximations of a task specification formula and encode the corresponding robustness degrees into the NLP cost functions [185, 160, 186, 187]. The work in [9] builds on the smooth approximation approach but focuses on handling multiple dynamic modes in robot manipulation tasks. [152, 188] apply the STL specifications and robustness degrees on the push recovery scenarios for bipedal locomotion.  [3] proposes a differentiable scheme for multi-arm manipulation problems by treating robot task assignment implicitly as continuous constraints that associate the states of robots and objects.
@@ VI-C Learning for combined TAMP
In recent years, there has been extensive exploration of learning techniques within TAMP community. Categorizing the works based on the roles and functionalities of the learned components, we primarily classify them into the following five categories. Note that, although some of the low-level motion planners used by works cited in this section might not be optimization-based, the methodologies discussed here are fundamental and highly relevant to optimization-based TAMP. A list of representative learning approaches is presented in table II.
@@
Table II: Overview of learning techniques for planning
@@ VI-C3 Learning Skill Policies
In recent advancements, some researchers have proposed the concept of acquiring foundational skills within the operational framework of TAMP systems. Notably these methods differ from conventional motion planning by learning skill policies, which are capable of generating trajectories through rollout and can be viewed as implicit motion plans. Learning skill policies at motion planning level has shown significant potential to improve the overall efficiency of TAMP due to the reusability and adaptability of the skills. McDonald et al. [33] propose to distill the knowledge of a TAMP system into a hierarchical RL policy, where they first leverage the TAMP solver to generate supervision data for imitation learning, and then the learned control policies are utilized to further speed up the TAMP solver. LEAGUE [32] proposes to learn RL policies with the guidance of a task planner, where the acquired skills are reused in the TAMP system to accelerate the learning of new skills, which progressively grows its capability for solving long-horizon manipulation tasks in a more efficient manner. The work in [204] extends [32] to leverage LLMs to guide skill learning. HITL-TAMP [191] develops an efficient teleoperation system that leverages TAMP to reach the beginning state of the demonstration phase, where the low-level control policies are learned with the collected data and integrated into the system in the testing stage. Meng et al. [205] propose to learn control policies for satisfying the long-term tasks specified with STL, where the learned models are used to generate trajectories via model predictive control during execution. Silver et al. [192] propose to jointly learn the symbolic operators and low-level skill policies with demonstration data, the extracted operators are first used to generate abstract task plans, and the learned policies are then invoked for generating motions to achieve subgoals.
@@ VII Future Challenges and Opportunities
Foundational Models for TAMP: The integration of LLMs and VLMs in robot planning is emerging but faces a few challenges. In robot planning, it is crucial to break down complex task specifications into actionable steps suited for particular environments [223, 224, 225, 63, 226]. However, LLMs and VLMs often struggle with this. Their plans may be too abstract, failing to consider the practical constraints of the physical world. This presents a significant issue for robots that need concrete and feasible instructions for physical operations. Furthermore, the current capabilities of LLMs and VLMs are limited, which in turn restricts their effectiveness in robot planning. For instance, spatial reasoning is essential in robot planning, yet LLMs and VLMs may not accurately understand physical spaces and dynamic environments [227, 228, 229]. Robot planning also requires considering historical data and long-term goals. The limited short-term memory of LLMs could result in information loss during sequential or multi-stage tasks, impacting the coherence and efficiency of planning [230, 231, 232]. Although recent developments aim to enhance the capabilities of LLMs and VLMs, they seem not to fundamentally solve these limitations [227, 233]. Additionally, in open-world settings such as homes, malls, and hospitals, robots also need the ability to adapt to new, unforeseen tasks. Despite the complexities, progress in developing LLMs and VLMs for these purposes is emerging [234, 235].
@@
Policy Learning in and for TAMP: In the realm of RL-based policy learning, key obstacles revolve around sample inefficiency and the reliance on meticulously crafted dense reward functions. These challenges often mandate the initial training of RL algorithms within simulators, thereby adding complexity to the transfer of acquired behaviors from simulation to real world [32]. While tapping into robot teleoperation data directly from real-world environments could alleviate generalization issues, this strategy demands careful system design to ensure seamless robot-human handover and maintain efficient data collection [191]. Moreover, the design of observation and action representations is crucial in enhancing the generalization and reusability of acquired skills.

### v4

@@ I Introduction
The main research focus in TAMP is to develop appropriate problem representations and algorithms that efficiently synthesize both symbolic and continuous components of the planning problem [2]. In the existing literature, there are three mainstream classes of TAMP methods: (i) constraint-based TAMP [4, 5], (ii) sampling-based TAMP [6, 7], and (iii) optimization-based TAMP [8, 9]. Constraint- and sampling-based TAMP characterizes the problem as a set of goal conditions. The solutions are typically found via constraint satisfaction or sampling-based approaches [10], which satisfy the defined goal conditions, but often cannot evaluate or compare the quality of the generated plan or the final state due to the lack of objective functions. In many robotics problems, goals are often expressed as an objective function rather than an explicitly defined set of states. For example, “Given a number of rectangular blocks on the table, build a stable structure that is as tall as possible with minimal robot control effort.” This is challenging for traditional sampling-based methods, which often require explicit goal definition and do not have mechanisms to compare plan qualities. As an exception, a specific class of sampling-based motion planning [11] have been proposed to address optimal planning using $\text{RRT}^{*}$ and $\text{PRM}^{*}$ [12, 13]. However, the complexity and expressiveness of the objective functions are often limited to simple costs such as path length, time, and energy consumption [14]. A comparison between optimization and sampling-based TAMP methods is presented in Table I. For clarification, the scope of this survey focuses on optimization-based TAMP, which naturally defines an objective function for representing the plan quality, in addition to task- and motion-level constraints. This framework enables us to represent and solve a broad range of tasks with complex objective functions.
@@
However, the hybrid optimization problems formed by optimization-based TAMP are often computationally intractable. A successful planning algorithm needs to simultaneously overcome the combinatorial complexity at the task planning level, and the numerical complexity at the motion planning level. As such, a common theme in optimization-based TAMP is to trade off between the complexity of the optimization and comprehensiveness of the information included in the planning problem. Either extreme of this trade-off tends to degrade either the quality or the computational efficiency of the resulting robot plans. Additionally, optimization-based TAMP faces several limitations comparing to sampling-based methods: (i) it is sensitive to the initial and goal conditions of the problem setup, which can lead to failures in complex environments such as complex obstacle geometry or difficult terrain, where certain initial and goal conditions can make it particularly challenging to find the optimal solution; (ii) the optimization results can be dependent on the initialization of decision variables, which might cause the planner to get stuck in local optima; (iii) optimization-based methods are not complete, meaning they cannot discover infeasible problems. Therefore, the challenge remains to improve the robustness of optimization-based TAMP and bridge the gap between planning for long-horizon tasks [22, 23] and generating highly dynamic robot behaviors, showcased in model-based optimal control strategies [24, 25, 26].
Aspect Optimization-Based TAMP Sampling-Based TAMP
Optimality Capable of converging to optimal or near-optimal solutions Asymptotically optimal but no guarantee on convergence speed
Task Definition Expressive definitions via objective functions & constraints Less expressive due to requirement for explicit goal definition, the objective function is limited to simple forms
Robot Dynamics Naturally incorporates dynamics in model-based TO Costly to handle dynamics via steering or forward propagation
Constraint Handling Handle complex constraints explicitly in model-based TO Rely on relaxation or simplification to handle manifold constraints
Robustness Sensitive to initialization and problem setup, potentially trapped by local optima Robust to variations in initial conditions due to random sampling
Completeness No mechanism to determine infeasibility Typically probabilistically complete
Table I: Comparison Between Optimization-Based vs Sampling-Based TAMP Methods
@@ I-A Survey Goals and Roadmap
This work is inspired by and builds upon previous surveys in TAMP [2, 10, 35, 36], but carries the unique overarching goal of reviewing the historical background and state-of-the-art optimization-based TAMP, and illustrating the connection between classical methods and the recent development in learning methods. Portions of this work are inspired by recent surveys in other relevant areas such as logic programming [23], formal methods [37], distributed optimization [38], and TO for legged locomotion [24]. Additionally, research contributions originated from 19 countries are highlighted to provide a global research landscape on TAMP innovations (Fig. 3).
Figure 3: Percentage statistics sorted by countries of origin. 182 references that directly address TAMP are included.
@@ IV-B Learning Models for Task Planning
More recently, deep learning techniques have been explored to learn the models from large-scale datasets. Ames et al. [83] propose to learn preconditions, action parameters, and effects from execution results of parameterized motor skills, which are then used to construct symbolic models for efficient planning. Neural task programming [84] proposes to learn neural models that recursively decompose a task demonstration video into robot executable action primitives. To further improve the generalization on long-horizon tasks, Neural task graphs [85] learns neural networks for generating conjugate task graphs, where the actions are represented as nodes and the dependencies between actions are modeled by edges, better exploring the compositionality. Regression planning networks [86] learns a neural model to iteratively predict the intermediate subgoals in a reverse order based on the current image observation and the final symbolic goal. Ceola et al. [87] propose to utilize deep reinforcement learning to train neural models for generating discrete actions. Deep affordance foresight [88] learns the long-term affordance of actions and the latent transition models to guide the search, and thereby informs the robot of the best actions to achieve the final task goal. Similarly, Liang et al. [89] propose to learn skill effect models that generate future terminal states of each parameterized skill, and then leverage these models to aid search-based task planning.
@@ IV-C LLMs for Task Planning
Traditionally, optimizing task plans for robots involves minimizing either the number of actions or the total plan cost, depending on whether action costs are considered. The emergence of LLMs, such as Google’s Bard, OpenAI’s ChatGPT [90], and Meta’s LLaMA [91], have reshaped the landscape of AI, including task planning for robots [92]. We categorize the LLM-based planning methods into the two groups: LLM-Native Planning Methods and LLM-Aided Planning Methods, where the former does not rely on external knowledge and the latter does, as shown in Fig. 8. Comparing to regular learning-based methods, LLMs are typically trained on a large amount of out-of-domain data that contains a great deal of commonsense knowledge. While LLMs are not strong in numerical reasoning (and hence optimization) [93, 94], the incorporation of LLMs improves the capabilities of natural language understanding, the acquisition of world knowledge, and commonsense reasoning. Such capabilities enable LLM-based planners to reason about symbolic information such as spatial relationships between objects [95] and symbolic correctness of a task sequence [30], without prior interaction with the robot environment. Therefore, LLMs as a task and domain agnostic reasoning module has the potential to enhance the scalability and generalizability of robot planning.
LLM-Native planning methods often incorporate additional components like reinforcement learning to enhance planning by choosing better actions. Conversely, LLM-Aided planning methods can be integrated with classical optimization strategies, ensuring satisfactory planning efficiency and practicality. These two approaches are compatible with optimization methods, while integrating LLMs enhances the overall planning capabilities.
@@ V-B Distributed Optimization
Spatial Structure: The spatial structure of the system can be exploited when the subsystems and their dynamics are separable. This property often exists in multi-robot systems, where the planning of each robot can be treated as a sub-problem. The decision variables are possibily coupled through the objective functions or the constraints (e.g. collision avoidance between robots). Local copies of the full state variables can be created for each robot to decouple the problem [119], as seen in Fig. 9(a). Readers are referred to [38, 120] for a detailed review of the multi-robot ADMM. Robustness is further studied in [121] given a multi-robot motion planning problem via ADMM. [122] accelerates TO for loco-manipulation tasks by modeling a quadruped robot with an articulated arm as three sub-robots.
@@ VI Integrated Task and Motion Planning
The optimization-based formulations for TAMP often involve a hybrid optimization of discrete symbolic-level decisions and continuous motion-level trajectories, as shown in Sec. II-B. To this end, we identified two general approaches to formulate and solve the hybrid optimization: logic-guided TAMP and TO-guided TAMP. While both approaches inherently solve the hybrid optimization problems, they fundamentally differ in formulations and algorithms, particularly in the definition of discrete variables and the selection of search spaces. Fig. 10 shows the overall organization of this section and main topics discussed. Table II presents a representative set of classical approaches for optimization-based TAMP, highlighting their formulations, algorithms, and whether dynamics or kinematics are considered in the application.
@@
Table II: Overview of classical approaches for optimization-based TAMP
@@ VI-A2 Multi-Modal Motion Planning Solved by TO
VI-A3 Receding Horizon TAMP
The real-world application of TAMP for long-horizon dynamic tasks is often hindered by failures in plan execution due to changes in the environment, interaction with human, or noisy sensor inputs. Receding horizon TAMP has been explored to mitigate this issue via online replanning. Receding horizon TAMP is analogous to model predictive control (MPC) [173], where the planning problem is solved iteratively over a receding time window. The specific challenge in receding horizon TAMP is to appropriately define the finite time horizon over the hybrid planning domain. The works in [174, 125] rely on task-specific decomposition, where each receding horizon planning iteration achieves the goal for a subtask. [175, 176] propose to plan over a fixed action-horizon, where a full task plan is generated in each iteration, while the motion plan is only computed for a predefined number of actions. [177] develops branch-MPC, where the objective function is optimized over a scenario tree, which is constructed by enumerating the predicted environmental responses.
@@ VI-B3 Smooth Approximation Methods
Different from the hierarchical and distributed methods above, another line of research to circumvent the combinatorial complexity of MIP is reformulate the hybrid optimization into a continuous NLP with a specific cost function or constraint representing the smooth approximation of the discrete task planning (Fig. 13(b)). Such a formulation can be solved more efficiently with gradient-based solvers. Recent works in STL utilize smooth approximations of a task specification formula and encode the corresponding robustness degrees into the NLP cost functions [191, 162, 192, 193]. The work in [9] builds on the smooth approximation approach but focuses on handling multiple dynamic modes in robot manipulation tasks. [154, 194] apply the STL specifications and robustness degrees on the push recovery scenarios for bipedal locomotion.  [3] proposes a differentiable scheme for multi-arm manipulation problems by treating robot task assignment implicitly as continuous constraints that associate the states of robots and objects. Analogous to receding horizon TAMP discussed in Sec. VI-A3, MPC-based methods have been developed based on STL to enable formal guarantees or reasoning about robustness of the task satisfaction in an online fashion [194, 195, 196, 197, 198].
@@ VI-C Learning for combined TAMP
Classical TAMP frameworks [2, 199] require accurate, special-purpose perception systems and hand engineered manipulation skills, rendering these approaches less effective while handling novel problems. To overcome this issue, in recent years, there has been extensive exploration of learning techniques within TAMP community. Data-driven approaches allow robots to make informed decisions based on prior examples and experiences, which enhance flexibility and generalizability. Furthermore, the scalability of classical TAMP methods is often limited by the problem size of the tree search for complex problems and the computational cost to evaluate heuristics and optimal trajectories. Learning-based methods show potential to accelerate or replace some of the computationally expensive components of classical methods, such as feasibility checking, search guidance, and skill learning. Categorizing the works based on the roles and functionalities of the learned components, we primarily classify them into the following five categories. Note that, although some of the low-level motion planners used by works cited in this section might not be optimization-based, the methodologies discussed here are fundamental and highly relevant to optimization-based TAMP. A list of representative learning approaches is presented in table III.
@@
Table III: Overview of learning techniques for planning
@@ VI-C2 Learning Search Guidance
For complex scenarios involving heterogeneous multi-agent systems, the planning framework must effectively handle task allocation and scheduling. Traditionally, these problems are addressed using heuristics-based approaches, but the combinatorial growth of the search space makes them computationally demanding. To tackle this challenge, reinforcement learning (RL) methods have been widely explored for both task allocation [215, 216] and task scheduling [217]. These studies demonstrate RL’s capability to significantly improve the time and memory requirements for multi-agent planning.
@@ VI-C3 Learning Skill Policies
In recent advancements, some researchers have proposed the concept of acquiring foundational skills within the operational framework of TAMP systems. Notably these methods differ from conventional motion planning by learning skill policies, which are capable of generating trajectories through rollout and can be viewed as implicit motion plans. Learning skill policies at motion planning level has shown significant potential to improve the overall efficiency of TAMP due to the reusability and adaptability of the skills. McDonald et al. [34] propose to distill the knowledge of a TAMP system into a hierarchical RL policy, where they first leverage the TAMP solver to generate supervision data for imitation learning, and then the learned control policies are utilized to further speed up the TAMP solver. LEAGUE [33] proposes to learn RL policies with the guidance of a task planner, where the acquired skills are reused in the TAMP system to accelerate the learning of new skills, which progressively grows its capability for solving long-horizon manipulation tasks in a more efficient manner. The work in [218] extends [33] to leverage LLMs to guide skill learning. HITL-TAMP [202] develops an efficient teleoperation system that leverages TAMP to reach the beginning state of the demonstration phase, where the low-level control policies are learned with the collected data and integrated into the system in the testing stage. Meng et al. [219] propose to learn control policies for satisfying the long-term tasks specified with STL, where the learned models are used to generate trajectories via MPC during execution. Silver et al. [203] propose to jointly learn the symbolic operators and low-level skill policies with demonstration data, the extracted operators are first used to generate abstract task plans, and the learned policies are then invoked for generating motions to achieve subgoals.
@@ VII Future Challenges and Opportunities
Foundation Models for TAMP: The integration of LLMs and VLMs in robot planning is emerging but faces a few challenges. In robot planning, it is crucial to break down complex task specifications into actionable steps suited for particular environments [237, 238, 239, 64, 240]. However, LLMs and VLMs often struggle with this. Their plans may be too abstract, failing to consider the practical constraints of the physical world. This presents a significant issue for robots that need concrete and feasible instructions for physical operations. Furthermore, the current capabilities of LLMs and VLMs are limited, which in turn restricts their effectiveness in robot planning. For instance, spatial reasoning is essential in robot planning, yet LLMs and VLMs may not accurately understand physical spaces and dynamic environments [241, 242, 243]. Robot planning also requires considering historical data and long-term goals. The limited short-term memory of LLMs could result in information loss during sequential or multi-stage tasks, impacting the coherence and efficiency of planning [244, 245, 246]. Although recent developments aim to enhance the capabilities of LLMs and VLMs, they seem not to fundamentally solve these limitations [241, 247]. Additionally, in open-world settings such as homes, malls, and hospitals, robots also need the ability to adapt to new, unforeseen tasks. Despite the complexities, progress in developing LLMs and VLMs for these purposes is emerging [248, 249].
Diffusion Models for TAMP: The use of diffusion models in motion planning, such as diffuser [250], decision diffuser [251] and diffusion policies [151, 252], has been explored for their flexibility and composability. [253] proposes a model-based diffusion planner that solves trajectory optimization using the diffusion process without external data. In the context of TAMP, diffusion models can serve as trajectory samplers for individual skills, offering a robust method for generating diverse and feasible motion plans. [254] presents generative skill chaining, where short-horizon skill-centric diffusion models are learned and a compositional framework is established to directly generate long-horizon plans given a plan skeleton. [255] integrates diffusion skill samplers into a classical TAMP method, which adapts the framework to partially observable planning domains. Promising future research can explore the synergy between LLMs and diffusion models to develop generative multi-task models and end-to-end TAMP frameworks.
Multi-Modal Sensing for TAMP: Currently, most TAMP frameworks rely predominantly on visual sensing. However, integrating multi-modal sensing, such as visual, tactile, and acoustic modules, can significantly improve the robots’ capabilities for contact-rich tasks in imitation learning [256, 257, 258], since each sensing modality provides unique and useful contact information related to the manipulation task that covers a wide range of geometric scales and frequency bandwidths. Future research opportunities include further integrating multi-modal sensing with TAMP by advanced sensor fusion as well as contact information representation, extraction, and utilization. Given these potentials, we can potentially enhance the performance of TAMP with better heuristics for task planning and more accurate contact models for trajectory optimization.
@@
Policy Learning in and for TAMP: In the realm of RL-based policy learning, key obstacles revolve around sample inefficiency (the cost of trial and error) and the reliance on meticulously crafted dense reward functions. These challenges are exacerbated in the realm of complex, long-term tasks and often mandate the initial training of RL algorithms within simulators, thereby adding complexity to the transfer of acquired behaviors from simulation to real world [33]. While tapping into robot teleoperation data directly from real-world environments could alleviate generalization issues, this strategy demands careful system design to ensure seamless robot-human handover and maintain efficient data collection [202]. Moreover, constructing diverse task configurations that encompass an adequate range of scenarios in terms of object geometries, spatial arrangements, and lighting conditions is nontrivial, presenting additional challenges for acquiring generalizable models in long-horizon tasks. Additionally, the design of observation and action representations is pivotal in enhancing the generalization and reusability of acquired skills.
@@
TAMP in Real-World Applications: Optimization-based TAMP has diverse real-world applications across various industries. In industrial settings, TAMP is used for construction planning [174], and rebar grid traversal [21]. Additionally, TAMP enables unmanned aerial vehicles (UAVs) to navigate complex environments for delivery services and environmental monitoring [271], as well as agricultural tasks [272]. In domestic applications, TAMP facilitates household tasks such as cooking [273], and manipulation of doors and dishwashers [19]. In lab environment, TAMP is deployed for medical test tube rearrangements [274]. Expanding TAMP to broader real-world applications requires overcoming current challenges such as developing more robust methods to accommodate varying environmental factors, problem settings, and human interactions. Additionally, proper scene understanding and representation that captures spatial and semantic relationships is essential and remains an open research area for deploying TAMP in open-world environments. Furthermore, implementing low-level control for robots in real-world scenarios involves handling environmental variability, sensor noise, and calibration issues, which affect feedback reliability. Addressing these challenges is crucial for effective and safe robot deployment.
Ethical and Societal Implications: The deployment of TAMP and robotics raise several ethical and societal considerations [275, 276], including (i) ensuring the safety and reliability of robot planning, particularly in human-populated environments; (ii) developing strategies to mitigate socio-economic impact of job displacement caused by increased automation; (iii) responsible handling of data collected during training of machine learning models to protect privacy and intellectual property; (iv) minimizing environmental impact via more energy efficient training process for machine learning models and sustainable practices in the the production, operation, and disposal of robots.
Glossaries of Terms
Alternating Direction Method of Multipliers (ADMM): An algorithm that solves convex optimization problems by breaking it into smaller pieces, each of which is easier to handle [100].
AI Planning: An area of artificial intelligence that studies the process of automated generation of sequence of actions to achieve specific goals [22].
Bilevel Optimization: A mathematical program, where an optimization problem contains another optimization problem as a constraint [277].
Foundation Model: A general purpose model, typically trained on a large amount of data, that can be adapted for various downstream applications [278].
Heuristics: Methods to organize the search space and guide the search algorithm [22].
Large Language Model: A large pretrained language model designed to understand and generate human-like text [92].
Mixed Integer Programming: An optimization problem with a combination of continuous and discrete decision variables [180].
Mixed Logical Dynamical System: A formulation of system dynamics with mixed integer constraints [37].
Model Predictive Control: A class of control methods where a model is used to predict the future of the controlled system over a receding planning horizon [173].
Motion Planning: The problem of moving a mechanical system from a start state to a goal region [11].
Nonlinear Programming: An optimization problem where the objective function or constraints are nonlinear [99].
Reinforcement Learning: A problem where an agent learns to make decisions through a goal-directed interaction with the uncertain environment [279].
Task and Motion Planning: A hybrid planning problem that integrates high-level task planning and low-level motion planning, which enables reasoning over long-horizon, dynamic tasks [2].
Task Planning: The problem of generating an action sequence that accomplishes the goal of the task and satisfies the task specifications [23].
Temporal Logic: The formal methods for describing time dependent rules and symbols, often used to specify the correctness of a finite-state transition system [37].
Trajectory Optimization: The problem of generating a continuous robot motion path and a control sequence that optimizes an objective function subject to a set of kinematics and/or dynamics constraints [99].